OpenStudy (anonymous):

how do i rewrite this expression using associative property of addition or multiplication ? 5/6(6/5m)

Parth (parthkohli):

5/6(6/5m)?? I don't understand what you're trying to comprehend.

OpenStudy (anonymous):

\[\LARGE \frac{5}{6}\left(\frac{6}{5}m\right)\] I don't know what you mean, but according to your post, this is what you've typed.

OpenStudy (anonymous):

so I guess this will help. http://www.aaastudy.com/mul74bx2.htm

OpenStudy (anonymous):

(2 * 3) * 4 = 2 * (3 * 4) Associative Property:

OpenStudy (anonymous):

yes kreshnik, that is how the problem is written , suppossed to rewrite it using associative property of addition or multiplication

OpenStudy (anonymous):

can you do it ? :)

Parth (parthkohli):

\[a + b + c = (a + b) + c = a + (b + c)\]

OpenStudy (anonymous):

having a hard time understanding the method

Parth (parthkohli):

@birdkeeper It's somewhat like the commutative property. You need not worry about the order in multiplication and addition. In a,b,c, you can add/multiply any two numbers first and continue.

OpenStudy (anonymous):

Example: \[\LARGE \frac{5}{2}\left({3\over 7}x\right)\] \[\LARGE \left(\frac{5}{2}\cdot {3\over 7}\right) x\]

Parth (parthkohli):

Let's have a real life numeral example: You have 3 apples. I give you 4 bananas. What do you have in total? Now I give you 2 more apples. Now, what number of fruits do you have in total? Another one, You have 3 apples. I give you 2 apples. What do you have? Now I give you 4 bananas. What do you have now? This one concentrates on the associative property. \[(3 + 4) + 2 = (3 + 2) + 4 = 9\]

OpenStudy (anonymous):

@ParthKohli Lol my math would suck according to bananas, I'm afraid I'd eat them all :( hahha.....

OpenStudy (anonymous):

your example makes it clearer to me now. thank you you so much

OpenStudy (anonymous):

@birdkeeper you like bananas too, don't you ? :P lol hahaha...

OpenStudy (anonymous):

yes i do like bananas. i would do the math then have pb and banana sandwich on wheat. Ha Ha

OpenStudy (anonymous):

hahahaahha...

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

yw